### v3

@@ 2.2 Coming-Out Event
Every year in June, LGBT Pride Month, a dedicated time to honor and celebrate the LGBT community, takes place. Since $2018$, Riot Games has actively participated in this month-long celebration by integrating new content into League of Legends during the month of June. This includes the introduction of in-game cosmetics, such as character skins, as well as emotes that allow players to express themselves in the game. It is important to note that while these additions enhance the visual and expressive elements of the game, they do not alter the game’s mechanics or the characteristics and abilities of the League of Legends characters.121212 We check this in Section 5.1, where we demonstrate that characters’ strength was unaffected by LGBT Pride Month.
Riot has also supported the representation of the LGBT community in the game by unveiling the sexual minority status of specific characters. In our paper, we primarily focus on the character Graves, chosen due to the existence of a well-defined announcement regarding his sexual orientation.131313 For instance, in $2018$, the character Neeko was confirmed as lesbian, marking her as the first openly LGBT character in the game. Furthermore, in $2021$, the characters Diana and Leona were revealed to be lesbians, while the character Nami was declared to be lesbian and polyamorous.
At the beginning of the $2022$ LGBT Pride Month, Riot Games released a short story featuring Graves and Twisted Fate. This story officially unveils Graves’ sexual orientation, establishing him as a gay character.141414 The story unveiling Graves’ sexual orientation also subtly hints at Twisted Fate’s pansexuality, although this is not explicitly stated. We investigate whether this implied revelation has captured the players’ attention in Appendix C. We highlight the relatively low attention directed towards Twisted Fate from players, who were primarily focused on Graves and the explicit establishment of his sexual orientation. As a result, we concentrate our analysis on Graves and his disclosure for a more credible identification of the effects of coming out. The following quotes provide two pivotal passages of the narrative:151515 The whole story is available at https://universe.leagueoflegends.com/en_SG/story/the-boys-and-bombolini/.
@@
This coming-out event closely approximates an ideal experiment where individuals randomly disclose their sexual minority status, thus providing a unique setting to identify the effects of coming out on players’ preferences for Graves.161616 It is crucial to distinguish between the coming-out event and the disclosure of Graves’ sexual orientation. The coming-out event encompasses both Graves’ disclosure and the start of LGBT Pride Month. While this is not a concern for identification, it requires careful interpretation of the findings. To maintain clarity, we generally refer to the effects of the coming-out event in our analysis. Further discussion on this topic is deferred to Section 5.5 and Appendix D.
@@
3 Data and Methodology
In this section, we introduce the data and explain the methodology employed in our analysis. The next subsection outlines the construction of our data sets and details how we gauge players’ revealed preferences for characters. We then provide a formal review of the synthetic control estimators employed to isolate the effects of the coming-out event on players’ preferences for Graves.
3.1 Data
We obtain our data by accessing the Riot Games API, which provides us with valuable information about League of Legends matches.
@@ 3.1 Data
The game operates on multiple servers located worldwide. Within these servers, we target the top tier of the League of Legends ranked system. By targeting this specific group of players, we aim to minimize the noise that may arise from players who are not fully engaged in the game, thus reducing the risk of attenuation bias.171717 Players can choose between draft or ranked matches. Both game modes share the same mechanics and objectives. However, while draft matches are more casual and do not have consequences for players’ rankings or ratings, in ranked matches players earn or lose points based on the outcome of the match to determine their position within the ranked system. By focusing on ranked matches we aim to further reduce the risk of attenuation bias. Furthermore, this focus increases the chances that players in our data set are aware of the coming-out event, thereby strengthening the credibility of our identification strategy.181818 We acknowledge that some of these players may have followers or engage in streaming activities, which could reduce the level of anonymity within our sample. Essentially, our focus trades off a degree of anonymity for reduced attenuation bias and strengthened credibility of identification.
Specifically, we identify the players within the top tier of each server as of July 2022. For each of these players, we gather raw data on the full history of matches they engaged in during the period January-July $2022$. This approach inherently includes, to a large extent, players who were not in the top tier as of July $2022$ but have been matched with our focal players during the time span we consider.
@@
To clean the raw data, we apply a multi-step procedure. First, we aggregate server-specific regions into four primary macro-regions—Latin America, North America, Europe, and Korea—and exclude servers where match coverage is sparse, namely Japan, Oceania, Russia, and Turkey.191919 The Latin America macro-region includes servers from North Latin America, South Latin America, and Brazil. Europe combines the North/East Europe and West/South Europe servers. Korea and North America are standalone servers. Further information on server structure and geographic coverage is available at https://wiki.leagueoflegends.com/en-us/Servers. Second, we eliminate duplicate observations, that arise when the same match appears multiple times due to its inclusion in the histories of multiple players. We also drop corrupted matches, defined as those with fewer than ten unique participants, a duration exceeding two hours, or missing data. Finally, we assign each character “main” and “auxiliary” roles based on historical usage patterns and exclude matches in which characters were used in atypical roles.202020 Characters in League of Legends are explicitly designed to fulfill a primary and a secondary role. We use observed historical play patterns to empirically assign each character their main and auxiliary roles. Excluding matches where characters are used outside these roles ensures that our measure of character usage reflects meaningful player choices, rather than experimental or off-meta behavior that could introduce noise into the analysis. After these steps, our clean sample comprises 136,399136{,}399 unique players participating in 146,451146{,}451 matches.
@@
We further construct an additional player-level daily panel that captures individual behavior over time. To ensure meaningful variation, we restrict the sample to players who participated in at least 10 matches both before and after the coming-out event, resulting in 7,4217{,}421 unique players. For each player, we compute the number of matches played and hours spent in-game each day, along with their daily win rates.212121 Hours spent in-game capture only the duration of matches. This substantially underestimates total time spent on the game, as it excludes time devoted to matchmaking, character selection, and other pre- and post-match activities. We also record the frequency with which players select specific characters and roles. This data set enables us to examine whether individual behavior adjusts in response to the coming-out event.
@@
Table 1 presents descriptive statistics on Graves usage and player behavior before and after the coming-out event. Graves’ pick rate declines from $18.56\%$ to $12.32\%$—a drop of more than $30\%$—suggesting a substantial shift in player preference for Graves.222222 Table A.I in Appendix A shows that, conditional on Graves’s primary role, it ranks as the second most frequently selected character before the disclosure. However, after the treatment, its ranking drops to fifth place. This decline is not accompanied by a deterioration in Graves’ win rate, which remains essentially flat. We further see no evidence that players disengaged from the game following Graves’ disclosure. On the contrary, both the average number of daily matches and hours spent in-game slightly increase post-treatment, ruling out attrition or general disengagement.232323 Figure A.I in Appendix A illustrates the distribution of players’ total and average daily activity, measured in both number of matches played and hours spent in-game. Finally, players’ role selection patterns remain largely stable across the pre- and post-treatment periods. The lack of major shifts suggests that players who stopped using Graves likely substituted toward other characters within the same role (jungle). We investigate the possibility of intra-role spillovers in greater detail in our robustness checks and find that our main results are unaffected.
3.2 Methodology

### v4

@@ 2.2 Coming-Out Event
Every year in June, LGBT Pride Month, a dedicated time to honor and celebrate the LGBT community, takes place. Since $2018$, Riot Games has actively participated in this month-long celebration by integrating new content into League of Legends during the month of June. This includes the introduction of in-game cosmetics, such as character skins, as well as emotes that allow players to express themselves in the game. It is important to note that while these additions enhance the visual and expressive elements of the game, they do not alter the game’s mechanics or the characteristics and abilities of the League of Legends characters.151515 We check this in Section 5.1, where we demonstrate that characters’ strength was unaffected by LGBT Pride Month.
Riot has also supported the representation of the LGBT community in the game by unveiling the sexual minority status of specific characters.161616 For instance, in $2018$, the character Neeko was confirmed as lesbian, marking her as the first openly LGBT character in the game. Furthermore, in $2021$, the characters Diana and Leona were revealed to be lesbians, while the character Nami was declared to be lesbian and polyamorous. In our paper, we primarily focus on the character Graves, chosen due to the existence of a well-defined announcement regarding his sexual orientation.
At the beginning of the $2022$ LGBT Pride Month, Riot Games released a short story featuring Graves and Twisted Fate. This story officially unveils Graves’ sexual orientation, establishing him as a gay character.171717 The story unveiling Graves’ sexual orientation also subtly hints at Twisted Fate’s pansexuality, although this is not explicitly stated. We investigate whether this implied revelation has captured the players’ attention in Supplementary Information E. We highlight the relatively low attention directed towards Twisted Fate from players, who were primarily focused on Graves and the explicit establishment of his sexual orientation. As a result, we concentrate our analysis on Graves and his disclosure for a more credible identification of the effects of coming out. The following quotes provide two pivotal passages of the narrative:181818 The short story was published on Riot Games’ official League of Legends Universe website and promoted as part of the 2022 Pride Month content campaign, making it an official element of the game’s narrative canon. The whole story is available at https://universe.leagueoflegends.com/en_SG/story/the-boys-and-bombolini/.
@@
This coming-out event closely approximates an ideal experiment where individuals randomly disclose their sexual minority status, thus providing a unique setting to identify the effects of coming out on players’ preferences for Graves.191919 It is crucial to distinguish between the coming-out event and the disclosure of Graves’ sexual orientation. The coming-out event encompasses both Graves’ disclosure and the start of LGBT Pride Month. While this is not a concern for identification, it requires careful interpretation of the findings. To maintain clarity, we generally refer to the effects of the coming-out event in our analysis. Further discussion on this topic is deferred to Section 5.5 and Supplementary Information F.
@@
3 Data
@@
We obtain our data by accessing the Riot Games API, which provides us with valuable information about League of Legends matches.202020 The API is publicly available and can be accessed using a standard developer key issued by Riot Games, allowing data collection without relying on web scraping. The game operates on multiple servers located worldwide. Within each server, we focus on players in the highest tier of the ranked system. By targeting this specific group of players, we aim to minimize behavioral noise arising from infrequent or weakly engaged players who may be less exposed to in-game information and whose character choices are more idiosyncratic, thereby reducing the risk of attenuation bias.212121 Players can choose between draft or ranked matches. Both game modes share the same mechanics and objectives. However, while draft matches are more casual and do not have consequences for players’ rankings or ratings, in ranked matches players earn or lose points based on the outcome of the match to determine their position within the ranked system. By focusing on ranked matches we aim to further reduce the risk of attenuation bias. Furthermore, this focus increases the chances that players in our data set are aware of the coming-out event, thereby strengthening the credibility of our identification strategy.
Specifically, we identify the players within the top tier of each server as of July 2022. For each of these players, we gather raw data on the full history of matches they engaged in during the period January-July 20222022.222222 A potential concern is that our sampling frame might exclude players who exited the game immediately following the disclosure. Several features of our data mitigate this concern. First, because we collect full match histories, the data set also includes all teammates and opponents appearing in those matches—many of whom are not part of the July top-tier population. This substantially broadens the effective sample and further reduces concerns that post-treatment exit removes players from observation. Second, as shown in Figure A.I in Appendix A, daily active player counts exhibit no decline around the disclosure date, either for the full sample or when distinguishing between prior Graves users and non-prior users. If anything, engagement remains stable or increases during the post-treatment period, consistent with the descriptive statistics reported below. This approach inherently includes, to a large extent, players who were not in the top tier as of July $2022$ but have been matched with our focal players during the time span we consider.232323 Our data are fully anonymized and do not contain demographic information. As a result, we cannot directly observe or report the gender, age, or other personal characteristics of the players in our estimation sample.
@@
To clean the raw data, we apply a multi-step procedure. First, we aggregate server-specific regions into four primary macro-regions—Latin America, North America, Europe, and Korea—and exclude servers where match coverage is sparse, namely Japan, Oceania, Russia, and Turkey.242424 The Latin America macro-region includes servers from North Latin America, South Latin America, and Brazil. Europe combines the North/East Europe and West/South Europe servers. Korea and North America are standalone servers. Further information on server structure and geographic coverage is available at https://wiki.leagueoflegends.com/en-us/Servers. Second, we eliminate duplicate observations, that arise when the same match appears multiple times due to its inclusion in the histories of multiple players. We also drop corrupted matches, defined as those with fewer than ten unique participants, a duration exceeding two hours, or missing data. Finally, we assign each character “main” and “auxiliary” roles based on historical usage patterns and exclude matches in which characters were used in atypical roles.252525 Characters in League of Legends are explicitly designed to fulfill a primary and a secondary role. We use observed historical play patterns to empirically assign each character their main and auxiliary roles. Excluding matches where characters are used outside these roles ensures that our measure of character usage reflects meaningful player choices, rather than experimental or off-meta behavior that could introduce noise into the analysis. After these steps, our clean sample comprises 136,399136{,}399 unique players participating in 146,451146{,}451 matches.
@@
We further construct an additional player-level daily panel that captures individual behavior over time. To ensure meaningful variation, we restrict the sample to players who participated in at least 10 matches both before and after the coming-out event, resulting in 7,4217{,}421 unique players. For each player, we compute the number of matches played and hours spent in-game each day, along with their daily win rates.262626 Hours spent in-game capture only the duration of matches. This substantially underestimates total time spent on the game, as it excludes time devoted to matchmaking, character selection, and other pre- and post-match activities. We also record the frequency with which players select specific characters and roles. This data set enables us to examine whether individual behavior adjusts in response to the coming-out event.
@@
Table 1 presents descriptive statistics on Graves usage and player behavior before and after the coming-out event. Graves’ pick rate declines from $18.56\%$ to $12.32\%$—a drop of more than $30\%$—suggesting a substantial shift in player preference for Graves.272727 Table A.I in Supplementary Information A shows that, conditional on Graves’s primary role, it ranks as the second most frequently selected character before the disclosure. However, after the treatment, its ranking drops to fifth place. This decline is not accompanied by a deterioration in Graves’ win rate, which remains essentially flat. We further see no evidence that players disengaged from the game following Graves’ disclosure. On the contrary, both the average number of daily matches and hours spent in-game slightly increase post-treatment, ruling out attrition or general disengagement.282828 Figure A.II in Supplementary Information A illustrates the distribution of players’ total and average daily activity, measured in both number of matches played and hours spent in-game. Finally, players’ role selection patterns remain largely stable across the pre- and post-treatment periods. The lack of major shifts suggests that players who stopped using Graves likely substituted toward other characters within the same role (jungle). We investigate the possibility of intra-role spillovers in greater detail in our robustness checks and find that our main results are unaffected.
4 Results
In this section, we present our main results. First, we present our main findings. We then investigate heterogeneity across players with different levels of in-game performance. Finally, we explore the possibility of regional variations in attitudes toward the LGB community by replicating our analysis across different macro-regions. Supplementary Information B provides a formal review of the synthetic control estimators employed to isolate the effects of the coming-out event on players’ preferences for Graves, while Supplementary Information C reports a series of robustness checks that support the reliability of our estimates.
4.1 Main Results
A simple comparison of Graves’ pick rates before and after the disclosure may not accurately reflect the impact of the coming-out event on players’ preferences for that character, as other factors could have changed during that period. To address this issue, we construct a synthetic control unit [[, see, e.g.,]]abadie2003economic, abadie2010synthetic, abadie2015comparative, abadie2021using, abadie2022synthetic by weighting other characters to approximate the pick rates of Graves before the disclosure. This method allows us to isolate the effects of the coming-out event on players’ revealed preferences for Graves and gain insight into how these preferences would have behaved in the absence of the disclosure. To mitigate the potential for spillover effects, we exclude Twisted Fate and other four characters (Diana, Leona, Nami, and Neeko) that were already members of the LGB community prior to the coming-out event from the donor pool.292929 Nevertheless, even if included in the donor pool, the estimator assigns these characters zero weight.
Figure 2: Graves’ daily pick rates and synthetic control estimation results. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 2 displays Graves’ actual and synthetic pick rate series.303030 Figure A.III in Supplementary Information A displays the identities and the contributions of the characters in the donor pool with non-zero estimated weights. Figure A.IV plots the corresponding pick rate series. Overall, our analysis suggests a substantial negative impact of the coming-out event on players’ preferences for Graves. Before the disclosure, the synthetic control estimator closely approximates the trajectory of Graves’ pick rates, providing support for the estimator’s ability to predict the counterfactual series. However, starting from June $1^{st}$, $2022$, the two series diverge substantially, with Graves’ pick rates consistently dropping below those of the synthetic control. This gap persists over time, extending even beyond the conclusion of LGBT Pride Month. The estimated average effect, reported in the first column of Table C.I in Supplementary Information C, is 7.11-7.11 percentage points and is statistically different from zero at the $5\%$ significance level.This implies a decline of $38.32\%$ of the pre-treatment average preferences for Graves.
We conduct a series of robustness checks, detailed in Supplementary Information C, all of which reinforce our main finding. First, our results are robust to alternative estimation strategies and donor pool compositions: we obtain consistent estimates using a regularized synthetic control estimator and when restricting the donor pool to characters unlikely to act as substitutes for Graves. This restriction directly addresses potential spillovers arising from substitution behavior: if players who stop selecting Graves mechanically substitute toward closely related characters, including those characters in the donor pool would violate a no-interference assumption by transmitting treatment effects into the synthetic counterfactual. Second, we assess predictive accuracy by artificially backdating the treatment [[, see, e.g.,]]abadie2022synthetic; discrepancies between the actual and synthetic series emerge only on the true disclosure date and closely match those in the main analysis, demonstrating the estimator’s ability to forecast the counterfactual series and supporting a no-anticipation assumption. Third, a leave-one-out analysis [[, see, e.g.,]]abadie2021using confirms that no single donor character drives the results. Fourth, we implement a within-role placebo design by replicating the synthetic control analysis for the four most frequently selected jungle characters (excluding Graves) in the pre-treatment period. None of these placebo units exhibits a comparable and sustained post-disclosure divergence from its synthetic counterpart, and we do not observe a synchronized decline in jungle pick rates around June 1, 2022. Overall, these patterns indicate that the estimated effect for Graves is not driven by broader shocks to role preferences, but is instead character-specific.
4.2 Player Skills Heterogeneity
To examine whether responses to the coming-out event differ by player ability, we conduct a heterogeneity analysis based on a proxy for player skill derived from pre-treatment in-game performance. Specifically, we compute each player’s win rate prior to the treatment date, restricting the sample to the same 7,4217{,}421 users selected for our player-level data set to ensure the reliability of this metric. We then classify these players into above- and below-median skill groups based on this metric. Using this stratification, we replicate our main synthetic control analysis separately for each group. The estimation proceeds identically to our baseline specification of the previous subsection.
This stratification allows us to investigate whether the decline in Graves’ pick rate is concentrated among particular subpopulations—such as relatively lower-performing players within the top-tier ranked population—or whether it reflects a broader behavioral response. Consistent estimates across skill groups would strengthen the external validity of our findings, mitigating concerns that the observed response is driven by a narrow or non-representative segment of the player base.
Figure A.V in Supplementary Information A plots the distribution of overall win rates and total matches for players in each skill group. Although the win rate distributions overlap substantially, they differ in their means, with players classified as higher-skilled exhibiting slightly higher win rates on average. In contrast, the distribution of total matches is nearly identical across groups, indicating that conditioning on pre-treatment win rate does not mechanically segment players by engagement level. This is important for interpretation: it implies that our stratification holds constant (approximately) overall game activity while isolating variation in a meaningful proxy for player quality.
(a) Players below median skill
(b) Players above median skill
Figure 3: Graves’ daily pick rates and synthetic control estimation results by skill group. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 3 presents the results. We observe a sustained decline in Graves’ pick rate immediately following the coming-out event in both the below- and above-median skill groups. The timing and persistence of the divergence from the synthetic counterfactual are remarkably similar across the two groups and closely mirror the patterns documented in our pooled specification.313131 Supplementary Information C shows that these findings are robust to backdating and leave-one-out exercises. These findings indicate that the estimated effect is not concentrated among players of a particular ability level but instead reflects a more widespread behavioral response within the top-tier ranked population.323232 While qualitatively similar, the pre-treatment pick rates of Graves are lower in both skill groups relative to the main specification. This difference arises from the additional sample restriction imposed here, whereby we include only players who participated in at least 10 matches both before and after the coming-out event to ensure a meaningful proxy for skill. The main specification, by contrast, includes the full player base, potentially encompassing more casual users with higher Graves usage. As such, differences in baseline levels should be interpreted as compositional.
4.3 Regional Heterogeneity
Previous research has documented substantial cross-country variation in social norms, legal environments, and the prevalence of individuals openly identifying as LGB [[, see, e.g.,]]badgett2020economic, badgett2021lgbtq.333333 An OECD report shows that even within a set of relatively comparable OECD countries, the size of the LGB communities differs by a factor of four [38]. To explore regional heterogeneity in behavioral responses to sexual minority disclosure, we exploit the geographic information associated with each match. Specifically, we divide the sample according to the macro-regions defined in Section 3 and apply our synthetic control methodologies separately within each region.
Figure 4: Graves’ daily pick rates and synthetic control estimation results by region. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month. Figure A.VI in Supplementary Information A displays the identities and the contributions of the characters in the donor pool with non-zero estimated weights.
Figure 4 displays the results. The synthetic control estimator closely approximates the trajectory of Graves’ pick rates for matches in Europe and Latin America before the disclosure, exhibiting pre-treatment root mean squared errors comparable to those of the pooled specifications. However, discrepancies arise in Korean and North American matches, where the pre-treatment root mean squared error is two to three times higher than that achieved with European and Latin American matches. In Europe, Korea, and Latin America, we estimate negative and persistent effects of the coming-out event on players’ preferences for Graves. Point estimates for the average treatment effect in these regions, reported in Table A.II in Supplementary Information A, are consistently negative and are statistically significant at the 5% level in our main specification. In North America, where the pre-treatment root mean squared is relatively large, we are unable to determine the sign of the impact, as the confidence intervals for the estimated average effect consistently encompass zero.
The estimated average effect varies substantially across regions. In our main specification, Europe exhibits the largest relative decline (-57.80% relative to pre-treatment average preferences), while Korea and Latin America show smaller declines of roughly -32%, indicating heterogeneous behavioral responses to the disclosure event. However, differences in magnitude should not be interpreted as direct measures of underlying social attitudes. Servers may differ along several dimensions that affect character selection independently of preferences toward sexual minorities. For example, regional differences in competitiveness could influence the degree of subjectivity in character choice. In addition, information about the disclosure may have diffused unevenly across regions.343434 Another factor may relate to how information about the disclosure diffused across regions. Although the announcement was officially localized into multiple in-game languages, players may also have learned about it through external channels such as social media, online forums, or streaming platforms, where exposure and salience can differ across linguistic and regional communities. We therefore carefully interpret these estimates as evidence of negative behavioral responses in several regions, rather than as a basis for comparing the intensity of underlying attitudes across regions.
5 Mechanisms
In Section 4, we established evidence of a substantial negative impact of the coming-out event on players’ revealed preferences for Graves. However, the players’ decision to switch from this character may reflect factors other than a response to the disclosure itself. This section discusses why such alternative explanations are unlikely, thereby strengthening the interpretation that the observed behavior reflects a response to the disclosure rather than confounding gameplay or design factors. Methodological details and supporting analyses are reported in Supplementary Information D.
5.1 Graves’ Strength
Essential for interpreting the estimated effects as responses to the disclosure is the fact that Graves’ strength remained unaffected by the coming-out event, as any change in character relative strengths could explain why players’ preferences shift away from Graves. To address this concern, we employ our synthetic control methodologies to examine the potential impact of the coming-out event on Graves’ strength. We measure characters’ strength using daily win rates.
As detailed in Supplementary Information D.1, our analysis reveals that the coming-out event had no impact on Graves’ strength, thus dismissing the possibility of a shift in his strength as an explanation for the results of Section 4. Moreover, players have real-time access to detailed statistics on each character’s performance from widely used online platforms, meaning they would have been aware that no abilities or attributes had changed during the period under study. These factors suggest that the negative impact of the coming-out event estimated in Section 4 is unlikely to be driven by actual or presumed changes in character relative strengths.
5.2 Engagement and Performance Across Usage Groups
If the observed decline in Graves’ usage were driven by gameplay considerations rather than the disclosure itself, we would expect systematic differences in engagement or performance across players with different pre-disclosure usage patterns.
To address this concern, we compare engagement and performance outcomes across players with different pre-treatment preferences for Graves. As detailed in Supplementary Information D.2, we classify players into prior users—those who selected Graves in at least $5\%$ of their pre-disclosure matches—and non-prior users, and compare engagement levels and performance before and after the coming-out event. Both groups play a similar number of daily matches, and neither shows signs of disengagement following the disclosure; if anything, match volumes increase slightly, suggesting players shifted to other characters rather than leaving the game. We also find no meaningful differences in win rates between or within groups, indicating that differences in pre-disclosure usage of Graves are not associated with differential performance or skill.
Overall, our findings provide no evidence that the post-disclosure decline in Graves’ usage is driven by gameplay considerations, thereby strengthening the interpretation that the observed behavior reflects a response to the disclosure rather than changes in in-game incentives.
5.3 Prior Usage and Performance
To ensure that our estimates capture behavioral responses to the disclosure rather than strategic considerations, it is crucial to assess whether shifting away from Graves to other characters impacts players’ performance. If there are performance costs, our estimates could be biased toward zero, as players might continue using Graves for strategic considerations. Moreover, if players switch characters primarily for convenience or strategic reasons, our analysis might capture a different phenomenon unrelated to the disclosure.
To investigate this possibility, we analyse whether players’ performance changed after moving away from Graves. As detailed in Supplementary Information D.3, we focus on prior Graves users and classify them into “treated” and “control” groups depending on the magnitude of their post-disclosure reduction in pick rate. We then employ difference-in-differences identification and estimation strategies [19] to compare the daily win rates of treated and control players before and after the disclosure.
Overall, we find that shifting away from Graves to other characters has no impact on players’ performance. None of the estimated effects is statistically different from zero, indicating that transitioning to other characters entails no performance-related costs. This reinforces the view that the decision to abandon Graves is unlikely to be driven by strategic or gameplay considerations.
5.4 New Substitute Character
On June $9^{th}$, $2022$, a new character (Bel’Veth) was released.353535 Figure A.X in Supplementary Information A displays Bel’Veth’s daily pick rates. Because Bel’Veth’s primary role overlaps with that of Graves, the new character can be considered a close substitute. Therefore, part of the post-June $9^{th}$ decline in Graves’ usage could, in principle, reflect players experimenting with the new character or seeking potential competitive advantages, rather than a response to his disclosure, thereby challenging our interpretation of the estimated effects.
If this substitution channel were the primary driver of the observed drop, we would expect a positive correlation between the magnitude of a player’s reduction in Graves usage and their subsequent adoption of Bel’Veth. However, the analysis in Supplementary Information D.4 shows that this is not the case. Players with substantial post-disclosure reductions in Graves usage are less likely to select Bel’Veth than those with moderate reductions, and even less likely than those who made no change at all. This pattern suggests that the release of Bel’Veth cannot account for the behavioral shift documented in Section 4.
5.5 Coming Out versus LGBT Pride Month
Supplementary Information F introduces a theoretical framework that formalizes the existence of two simultaneous treatments, discusses the interpretation challenges it poses, and outlines sufficient conditions for separating the effects of the coming out from those of Pride Month. Here, we provide the main intuitions behind our approach, directing the reader to the supplementary information for technical details.
To examine the potential impact of LGBT Pride Month on players’ preferences for LGB characters, we leverage the existence in our data set of other four characters already acknowledged as part of the LGB community before the coming-out event. These characters are subject only to a part of our treatment, specifically being part of the LGB community while LGBT Pride Month is ongoing, whereas Graves experiences both the disclosure of his sexual orientation and LGBT Pride Month. We create a composite LGB unit by averaging their daily pick rates and apply our synthetic control methodologies to measure the effect of Pride Month on players’ preferences for LGB characters. Under the assumption that Pride Month’s influence is homogeneous across all LGB characters, the difference between the Graves and composite LGB effects identifies the impact of his disclosure. Intuitively, if the estimated impact of LGBT Pride Month on players’ preferences for LGB characters is small relative to the estimated impact of the coming-out event on players’ preferences for Graves, this suggests that the findings of Section 4 must be primarily attributed to Graves’ disclosure.
The results, reported in Supplementary Information D.5, show that Pride Month had no measurable effect on players’ preferences for LGB characters. Under the homogeneity assumption discussed above, these findings support the interpretation that the estimated effects presented in Section 4 are primarily driven by Graves’ disclosure rather than being influenced by the broader context of LGBT Pride Month.
6 Conclusion
Disclosure of sexual orientation can have important social and economic implications. Understanding how individuals respond to such disclosure is therefore relevant for evaluating the broader consequences of coming out. Prior research documents that differential treatment and social exclusion of LGB individuals in educational, health, social, and political settings can lead to losses in human capital with detrimental economic effects [8], as well as severe long-term health consequences [13]. Examining behavioral responses to sexual orientation disclosure contributes to this broader discussion by providing evidence on how disclosure may shape subsequent interactions.
In this study, we utilize a comprehensive data set sourced from the widely popular online video game League of Legends and exploit exogenous variation in the identity of a playable character to credibly identify behavioral responses to sexual minority disclosure. Players in the game select a playable character before each match, and each character is characterized by game-relevant attributes and a background story. Leveraging an unexpected revelation during the 2022 LGBT Pride Month—wherein game developers disclosed the sexual orientation of a playable character—we examine how players adjust their revealed preferences following this disclosure. By tracking character selection over a meaningful period, we document a substantial and persistent decline in demand, with preferences decreasing by around $38\%$. These findings highlight that sexual orientation disclosure can meaningfully affect subsequent behavior in interactive environments and contribute to understanding how disclosure may shape social and economic interactions.
To strengthen the interpretation of the estimated effects as responses to the disclosure, we address and eliminate several alternative channels. First, we rule out the possibility that shifts in characters’ relative strengths could explain our estimated effect. Second, we show that players’ skills have no correlation with the choice to drop the character. Third, we provide evidence that switching to other characters does not affect the performance of the players involved and that the release of new characters in the post-treatment period is unlikely to serve as primary explanation behind the results. Fourth, we introduce a theoretical framework that formalizes the existence of two “simultaneous treatments” and use information on other playable characters that belong to the LGB group to rule out that LGBT Pride Month serves as the main explanation for the observed result.
While we do not distinguish between out-group animus or reduced identification and in-group preference, both mechanisms imply economically meaningful behavioral costs of identity disclosure. More broadly, the results highlight that identity revelation can shift demand in high-visibility environments in ways that persist beyond short-lived attention shocks, even absent material payoff changes. Our findings have broad implications for understanding how disclosure of sexual orientation can shape subsequent interactions and behavior. Raising awareness about behavioral responses to sexual orientation disclosure may contribute to designing interventions that promote inclusive and supportive communities [44].
Beyond societal implications, our findings are relevant for platform and content design. Then decline in Graves’ usage following the disclosure suggests that inclusive representation can generate immediate behavioral responses among some players. However, our design does not allow us to assess longer-run outcomes such as engagement, retention, brand value, or community norms, which are likely central to platform strategy.363636 Riot Games has continued to release official Pride Month content in 2023, 2024, and 2025, maintaining a public commitment to inclusive representation. From a welfare perspective, simmediate demand responses must be weighed against potential longer-run benefits of fostering inclusive environments and attracting more diverse user populations. Understanding how identity-based representation affects platform growth and long-term engagement remains an important direction for future research.
Supplementary Information A Additional Figures and Tables
Figure A.I: Daily number of active players. The solid black line shows the total number of active players, while the dashed red and dotted blue lines report activity separately for players with high and low pre-treatment Graves usage, respectively. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure A.II: Histograms of players’ total and average daily activity, measured in both number of matches played and hours spent in-game.
Figure A.III: Identities and contributions of characters in the donor pool for the Graves’ synthetic control displayed in Figure 2.
Figure A.IV: Daily pick rates of the characters contributing to the Graves’ synthetic control displayed in Figure 2. The rainbow area highlights LGBT Pride Month.
Figure A.V: Distribution of overall win rates and total matches played for players classified as above or below the median pre-treatment win rate.
Figure A.VI: Identities and contributions of characters in the donor pool for the Graves’ synthetic controls displayed in Figure 4.
Figure A.VII: Shares of matches played in each role before and after the coming-out event. Players are divided into two groups based on their preferences for Graves before his disclosure.
Figure A.VIII: Distribution of changes in average Graves’ pick rates (percentage points) between the pre- and post-disclosure periods among prior users. Negative values indicate increases in pick rates, while positive values indicate reductions.
Figure A.IX: Point estimates and simultaneous $95\%$ confidence bands allowing for clustering at the player level for the placebo $ATT(t)$. Each row corresponds to a different version of the treatment.
Figure A.X: Bel’Veth’s daily pick rates. The rainbow area highlights LGBT Pride Month.
Table A.I: Most popular characters.
Table A.II: Regional results.
Synthetic Controls Regularized Synthetic Controls (1) (2) (3) (4) All characters Only non-substitutes All characters Only non-substitutes Panel 1: Europe $\hat{\tau} -$8.416 -9.968 -6.819 -7.170 $p\text{-value}$ 0.004 0.021 0.008 0.120 $95\%$ CI [-14.102, -2.730] [-18.462, -1.475] [-11.890, -1.749] [-16.393, 2.040] N. Donors 7 7 17 15 RMSE 3.317 3.605 3.450 3.659 Pre-treatment average 14.62 14.62 14.62 14.62 Panel 2: Korea $\hat{\tau} -$10.637 -9.963 -9.461 -7.300 $p\text{-value}$ 0.003 0.086 0.011 0.250 $95\%$ CI [-17.649, -3.625] [-21.353, 1.428] [-16.791, -2.132] [-19.743, 5.138] N. Donors 3 2 5 4 RMSE 8.373 8.793 8.810 9.070 Pre-treatment average 31.409 31.409 31.409 31.409 Panel 3: Latin America $\hat{\tau} -$5.188 -5.225 -3.407 -3.200 $p\text{-value}$ 0.019 0.150 0.125 0.290 $95\%$ CI [ -9.517, -0.858] [-12.345, 1.895] [ -7.762, 0.948] [ -9.235, 2.820] N. Donors 5 5 12 11 RMSE 3.901 3.927 3.967 4.027 Pre-treatment average 16.229 16.229 16.229 16.229 Panel 4: North America $\hat{\tau}$ 0.597 0.597 1.587 1.830 $p\text{-value}$ 0.814 0.877 0.585 0.660 $95\%$ CI [ -4.382, 5.576] [ -6.961, 8.155] [ -4.109, 7.283] [ -6.395, 10.074] N. Donors 4 4 6 6 RMSE 6.237 6.237 6.351 6.382 Pre-treatment average 22.281 22.281 22.281 22.281
Notes. Point estimates, $p$-values, and $95\%$ confidence intervals for $\hat{\tau}$. Additionally, the number of donors receiving a non-zero weight and the pre-treatment root mean squared error are displayed. Each panel reports the results obtained using only matches from a particular macro-region. Each column corresponds to a different specification, with the specifications differing solely in the employed estimator and donor pool composition.
Supplementary Information B Methods